OpenStudy (anonymous):

Carla has 1 gallon of paint. She is going to pour it into a paint tray that measures 10 inches wide, 10 inches long, and 7 cm deep. (1 gallon = 231 in3, 1 inch = 2.54 cm) Which of the following scenarios will happen? The paint will not fill the tray by 44.59 in3. The paint will not fill the tray by 469 cm3. The paint will fill the tray exactly. The paint will overfill the tray by 44.59 in3. HELP PLZ

OpenStudy (anonymous):

paint tray volume = 10*10*(7/2.54)=275.59 in3 volume remain in tray after pouring paint = 275.59-231 =44.59 Hence paint will not fill tray by 44.59 in3

OpenStudy (anonymous):

thx i really appreciate